A company has established that the relationship between sale price for one of its product and the quantity sold per month is approximately D = 780 -10p units (D is the demand or quantity sold per month, and p is the price in pesos). The fixed cost is P800 per month, and the variable cost is P30 per unit produced. What number of units, D*, should be produced per month and sold to maximize net profit? What is the maximum profit per month related to the product?
1
Expert's answer
2020-09-09T11:02:24-0400

ANSWERS


D* = 240 units per month


Maximum profit (π) = P4,960 per month


SOLUTIONS

We are given "D = 780 - 10P \space units"

"Fixed \space costs (FC) = P800 \space per \space month"

"Variable \space costs = P30 \space per \space unit"


The inverse monthly demand function is therefore given by:

"10P = 780 - D"

"P = \dfrac {780}{10} - \dfrac {D}{10}"

"P = 78 - 0.10D"


"Price (P) = Average \space revenue (AR)"

"\therefore AR = 78 - 0.10D"


Total revenue (TR) is given by:

"TR = AR \u00d7 D"

"= (78 - 0.10D) \u00d7 D"

"= 78D - 0.10D^2"


Total cost = Variable costs + Fixed costs

"TC = VC + FC"

"= (P30 \u00d7 D) + P800"

"= 800 + 30D"


Total profit is given by:

Total profit = Total revenue - Total costs

"\u03c0 = TR - TC" "\u03c0 = (78D - 0.10D^2) - (30D + 800)"

"= 78D - 30D - 0.10D^2 - 800"

"= -800 + 48D - 0.10D^2"


Maximum profit is approached through differential calculus

"\dfrac {dπ}{dD} = \dfrac {d}{dD}(-800 + 48D - 0.10D^2)"

"= 48 - 0.20D"



"\dfrac {d^2π}{dD^2}= \dfrac {d^2}{dD^2}(48-0.20D)"


"= -0.20"


"Since \space \dfrac {d^2π}{dD^2} < 0" , profit(π) is maximum.


When profit (π) is maximum,

"\dfrac {dπ}{dD} = 0"


"=> 48 - 0.20D = 0"

"0.20D = 48"

"D = \dfrac {48}{0.20}"

"= 240 \space units"

"\therefore D^* = 240 \space units \space per \space month"


Calculating maximum profit (π)

"Max(\u03c0) = -800 + 48D^* - 0.10(D^*)^2"

"= -800 + 48(240) - 0.10(240^2)"

"= -800 + 11 520 - 5760"

"= P4,960"


"\therefore Max(π) = P4,960 \space per \space month"